Status Report

ISS On-Orbit Status 6 Nov 2002

By SpaceRef Editor
November 6, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously
or below. Hard work is the order of the day on the outpost in space.

For the visiting crew (VC-4), science activities continued with a packed
schedule, again starting the day off with the experiments VIRUS (collection
of saliva samples) and SLEEP (filling out a study questionnaire to evaluate
sleep performance/experience).

Main research activity today centered on the last session with CARDIOCOG
(study of the effects of microgravity on the cardiovascular system expressed
in the peripheral arteries, and the vegetative regulation of arterial
blood pressure and heart rate). [For each of the three subjects, Sergei
Zalyotin, Frank De Winne and Yuri Lonchakov, the Nikon F-5 and Nikon D1
still cameras and the DSR-PD150P video camera were set up to take general
and close-up imagery during experiment runs. The electrode-equipped subjects
took electrocardiogram (ECG) readings and arterial blood pressure measurements.
After final ops, the equipment was to be torn down by VC CDR Zalyotin,
and the hazardous thionyl chloride batteries were to be sealed and stored
safely in the Soyuz TM-34 OM (orbital module).]

VC FE-1 De Winne unstowed another sample for the MESSAGE experiment (effects
of micro-G on bacterial gene expression with special attention to genes
involved in the response to stress) and started its biological strain
incubation. [Incubated specimens are stored in the Russian Cryogem-03
refrigerator until the payload is dismantled by the VC-4 crew on departure
day (11/9).]

Zalyotin performed another temperature check (37 degC) of the running
AQUARIUS-B, then terminated the experiment and disassembled its equipment
for stowage.

Working on the MSG (microgravity science glovebox) in the Lab, Frank
De Winne removed tape #5 of the PROMISS experiment (protein crystal growth
monitoring by digital holographic microscope) and installed tape #6.

The third MSG experiment, NANOSLAB (study of aggregation mechanisms and
kinetics of ZSM-5 and Silicate-1 nanoslabs into hybrid phases under micro-G
conditions), failed last night after its activation. Troubleshooting was
initiated to recover the payload.

Yuri Lonchakov worked on the Diatomeya program of world ocean observations,
performing still photography with the Nikon-F5 camera during two observing
sessions. [Targets of interest included fields of organic and mineral
suspended matter in the dynamically active Atlantic waters off the coast
of Africa, as well as bioproductive waters in the East Equatorial Pacific
region. Lonchakov focused the Nikon-F5 on the mouth of the Cavalla River,
the mixing area of the cold and warm waters of the Guinea and Benguela
currents, the head of the South Equatorial Current, and the Pacific South
Equatorial Current.]

Yuri also took photos, with the Kodak 760 with 800mm focal-length lens,
for the Uragan program, targeting again the activity of the volcano El
Reventador in Ecuador. Images are later transferred from the camera storage
card to the EGE-2 computer hard-disk drive (HDD) for return to Earth.

The VC-4 crew was scheduled to conduct a new experiment program, EDUCATION
(EDC), consisting of video shooting of simple physical phenomena demonstrations
during orbital flight for education purposes. [De Winne and Lonchakov
were to use transparent drink bags, straws, a balsa wood plane and simple
items like pens, souvenirs and rubber seals to demonstrate the formation
of air bubbles in liquid, water drops on hard surface, flight of balsa
plane in zero-G, and throwing solid objects vs. letting them "fall".]

The ISS crew again had time set aside on their schedule for prepacking
(organizing equipment) and preparing for their departure on 11A.

FE-1/SO Peggy Whitson worked on EVA tool preparations. Her scheduled
cleanup of the Airlock Crewlock compartment had to be deferred since the
Russian Orlan-M suits are still in the Node. She was to use the time to
locate some missing equipment items.

Later in the day, Whitson and Korzun, supported by conference with the
ground, completed an MSS (mobile service system) refresher session to
hone their skills on the SSRMS (space station remote manipulator system).
[Peggy and Valery will operate the Canadarm2 for transfer of the P1 truss
on 11A (11/13), during which the truss element will be handed over from
the Shuttle RMS/Canadarm1, operated by STS-113 CDR Jim Wetherbee, to the

At 7:30am EST, the 24-hr. Holter cardiac activity monitoring ended for
CDR Valery Korzun. Before he doffed the ECG (electrocardiograph) harness,
he completed the periodic fitness evaluation (PFE). The PFE was also conducted
by Peggy Whitson, using the U.S. BP/ECG (blood pressure/electrocardiograph)
equipment plugged in a UOP (utility outlet panel).

Sergei Treschev and Yuri Lonchakov removed no-longer-needed electronic
equipment from the Soyuz TM-34/4S OM (orbital compartment), viz., a local
signal switch device (LKT/TA251) of the BITS onboard telemetry system,
with its ROM memory (PZU).

Troubleshooting of the CDRA (carbon dioxide removal assembly) continues
at high priority. The machine was successfully activated just prior to
crew sleep last night and ran for most of the sleep period, but not without
problems. It still continues to shut down due to the valve problem the
ground has been monitoring over the past few days. The decision was made
to transition the LSYS3 MDM back to the R1 software version, in case the
valve travel anomaly is somehow related to the recent loading of new software
on the Lab systems computers. [The LSYS MDMs were upgraded on 10/21 to
the R2 software version, and all three Lab systems computers were afterwards
functioning nominally.]

The crew was requested to verify an inventory assessment of available
LiOH (lithium hydroxide) canisters and their expiration dates.

From 6:35am to about 3:10pm EST, MCC-M (TsUP) performed fuel transfer
from Progress M1-9 (9P) to the ISS, first to the FGB tanks and then from
there to the SM tankage.

Later in the day, from 6:30pm to 11:00pm EST, TsUP will conduct another
solar array efficiency test of the SM and FGB photovoltaic cells, periodically
required to evaluate their performance over time. The test requires the
ISS to fly in XPOP attitude (X-axis perpendicular to orbit plane), which
is currently maintained.

During the solar array efficiency test, MCC-H, in coordination with TsUP,
will perform a calibration test on the RACU 6 (Russian-to-American converter
unit #6). [US and Russian specialists will compare power draws seen by
Houston and Moscow. These data will aid the development of appropriate
calibration coefficients for RACU power draw, necessary to accurately
coordinate Russian segment to US segment power transfers.]

At 3:20pm EST, the Expedition 5 crew were scheduled to participate in
a 20-min. televised downlink event (Ku- and S-band) with CNN (Miles O’Brien)
and possibly CNN’s "Talk Back Live" show.

More PAO events are scheduled for Frank De Winne and Yuri Lonchakov with
Belgian TV stations. Also, De Winne will have a private telecon via IP
(Internet protocol) phone and later will conduct his regular daily activities
downlink to Belgium.

CDR Korzun completed the daily routine maintenance of the SOSh life support
system, while Treschev prepared the IMS (inventory management system)
delta file for downlink, and FE-1/SO Peggy Whitson conducted the regular
status checkup of Lab payloads.

Today’s targets of the CEO (crew earth observations) program were Barcelona,
Spain (nadir pass. ESC [electronic still camera] requested), Khartoum,
Sudan (nadir pass over this city at the confluence of the Blue and White
Niles. ESC), W African fires (Dynamic event. Sahelian countries on the
bulge of Africa are experiencing a large number of fires. Looking nadir
and left), St. Thomas I., Caribbean (nadir pass. ESC), St. John I., Caribbean
(nadir pass. ESC), St. Croix, Caribbean (nadir pass. ESC), Salt Lake City,
Utah (nadir pass and a touch left. ESC), Albuquerque, New Mexico (nadir
pass and a touch right. ESC), Baja cloud patterns (Dynamic event. Numerous
vortices being shed from Baja’s coastal promontories), Necker reef, Hawaiian
chain (nadir pass. ESC), Nihoa reef, Hawaiian chain (nadir pass. ESC),
Congo-Zimbabwe Biomass Burning (looking left for large number of fires
in central-northern Mozambique. These fires are making news), and Patagonian
Glaciers (looking left of track).

CEO images can be viewed at the website:

U.S. and Russian Segment Status (as of 3:00 am EST).

Environmental Control and Life Support (ECLSS) and Thermal Control

Elektron O2 generator is powered On (48 amp mode; on backup pump). Vozdukh
CO2 scrubber is ON in AUTO air flow mode (ppCO2 limit: 4 mmHg). U.S. CDRA
CO2 scrubber is Off. TCCS is operational. BMP Harmful Impurities unit:
Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

Propulsion System (PS): Total propellant load available (SM + FGB + Progress)
— 3903 kg (8605 lb) [as of 10/24/02].

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and
BGA 4B in Autotrack mode (solar-tracking).
SM batteries: All batteries (8) in "Partial Charge" mode.
FGB batteries: Battery #1 is offline; all other batteries (5) are in "Partial
Charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is On; SKV-2 is Off.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
EXT-1 is On (primary), EXT-2 is off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line.
State vector — U.S. SIGI-1
Attitude — Russian segment (RS)
Angular rates — U.S. RGA-1 (from RS attitude)

Communications & Tracking Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF3 (mobile base system/power & data grapple
fixture 3) and 11A park position, with Keep Alive power on both strings.
MBS: Keep Alive power on both strings. POA: Keep Alive power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:37am EST [= epoch]):
Mean altitude — 392.9 km
Apogee — 402.6 km
Perigee — 383.1 km
Period — 92.41 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.001438
Orbits per 24-hr. day — 15.58
Solar Beta Angle — -23.0 deg (magnitude decreasing)
Altitude loss — 260 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ’98) — 22617
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane =
"sun-fixed" [yaw: -5.0, pitch: -8.6 deg., roll: 0.4 deg]),with
CMG/Thruster Assist Momentum Management).

For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.